OpenStudy (anonymous):

evaluate lim (x->0) of (sinx-x)/x^3

OpenStudy (dumbcow):

apply L'Hopitals rule differentiate \[\rightarrow \lim_{? \rightarrow ?} \frac{\cos x -1}{3x^{2}}\] still indeterminate so repeat and differentiate again \[\rightarrow \lim_{? \rightarrow ?} \frac{-\sin x }{6x}\] still indeterminate so repeat and differentiate again \[\rightarrow \lim_{? \rightarrow ?} \frac{-\cos x }{6} = -\frac{1}{6}\]

OpenStudy (zarkon):

If you know the taylor expansion of sine you can do the following \[\lim_{x\to0}\frac{\sin(x)-x}{x^3}=\lim_{x\to0}\frac{x-\frac{x^3}{6}+O(x^5)-x}{x^3}\] \[=\lim_{x\to0}\frac{-\frac{x^3}{6}+O(x^5)}{x^3}=-\frac{1}{6}+0=-\frac{1}{6}\]